OpenStudy (anonymous):

please i really need help!!! Use inverse operations to solve. Explain each step. 19 =h/3 – 8

OpenStudy (anonymous):

try to isolate h alone first + 8 both sides then times 3 both sides

OpenStudy (anonymous):

i still dont get it?

OpenStudy (anonymous):

\[19 = \frac{h}{3} - 8\]first we move the 8 to the other side by + 8 both sides\[19+8 = \frac{h}{3}-8 + 8\]\[27 = \frac{h}3 \]then to get rid of the 3, we times both sides by 3\[27 \times 3= \frac{h \times 3}{3}\]\[81 = h\]